# Nonlinear Ringdowns as Sources and Detectors of Quantum Gravitational Waves

### v1

@@ C.2 Multimode vacuum corrections and few-mode Hamiltonian
The evolution operator (46) contains contributions from all modes of the gravitational field, most of which are in the vacuum state. Following [13], we now show that the contribution due to empty modes can be safely neglected, as expected. The evolution of the cavity annihilation operator is given by,
$c(t)$ $=$ $\left(\prod_{k}U_{k}(t)\right)^{\dagger}c\left(\prod_{k}U_{k}(t)\right)$ (51)
@@
$|0\rangle=\prod_{k}|0_{k}\rangle$ (52)
@@
$c(t)=e^{i\mathcal{F}(t)c^{\dagger}c}e^{i\mathcal{F}(t)/2}\mathcal{G}(t)\ c$ (53)
@@
$\mathcal{G}(t)=\prod_{k}\langle 0_{k}|D(q_{k}\gamma_{k})|0_{k}\rangle=\exp\left[-\dfrac{1}{2}\sum_{k}q_{k}^{2}|\gamma_{k}|^{2}\right]$ (55)
@@
$\mathcal{F}(t)=\int\dfrac{d^{3}\bm{k}}{\sqrt{(2\pi)^{3}}}2\omega_{c}^{2}\left(\dfrac{8\pi G}{k^{3}}\right)\left(\omega_{k}t-\sin\omega_{k}t\right)\approx\left(\dfrac{\omega_{c}}{E_{\rm pl}}\right)^{2}\left(\int_{E_{\rm IR}}^{E_{\rm pl}}dk\right)t\approx\left(\dfrac{\omega_{c}}{E_{\rm pl}}\right)\omega t$ (56)
@@
$\mathcal{G}(t)=\exp\left(-\dfrac{1}{2}\int\dfrac{d^{3}\bm{k}}{\sqrt{(2\pi)^{3}}}\omega_{c}^{2}\left(\dfrac{8\pi G}{k^{3}}\right)|\gamma_{k}|^{2}\right)\approx\exp\left(-2\left(\dfrac{\omega_{c}}{E_{\rm pl}}\right)^{2}\int_{E_{\rm IR}}^{E_{\rm pl}}\dfrac{dk}{k}\right)=\exp\left(-2\left(\dfrac{\omega_{c}}{E_{\rm pl}}\right)^{2}\ln\left(\dfrac{E_{\rm pl}}{E_{\rm IR}}\right)\right)$ (57)
@@
Instead of the vacuum, we could consider all modes to be populated by a thermal state at about $1\text{\,}\mathrm{K}$, the expected temperature for the cosmic GW background [87]. This would not alter $\mathcal{F}(t)$, which is state-independent, so the estimates in (56) remain. The $\mathcal{G}(t)$ term, however, would acquire a correction factor at most $e^{-q_{\rm pl}^{2}(2\bar{n}+1)}$, with $\bar{n}=1/(e^{\hbar\Omega_{k}/k_{B}T}-1)$ and $q_{\rm pl}$ the GW coupling strength at the Planck frequency [13]. For a GW mode of $10\text{\,}\mathrm{H}\mathrm{z}$, peak of the expected cosmic GW background spectrum, this correction factor amounts to $\approx e^{-10^{-47}}$.
@@ C.3 Few mode Hamiltonian
$H_{O}=\omega_{c}c^{\dagger}c+\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-c^{\dagger}c\left(g_{a}Q_{a}+g_{b}Q_{b}\right)$ (58)
@@
$P_{a}=i(a^{\dagger}-a)\ ,\ P_{b}=i(b^{\dagger}-b)$ (60)
in accordance do Eq. (73). Note the two modes couple differently to the GW detector; we have,
$\frac{q_{a}}{q_{b}}=\left(\frac{\omega_{b}}{\omega_{a}}\right)^{3/2}=2\sqrt{2}$ (61)
@@
$U(t)=e^{-iB(t)(c^{\dagger}c)^{2}}e^{q_{a}c^{\dagger}c(\gamma_{a}a^{\dagger}-\gamma_{a}^{*}a)}e^{q_{b}c^{\dagger}c(\gamma_{b}b^{\dagger}-\gamma_{b}^{*}b)}$ (62)
with $B(t)=\sum_{k=a,b}B_{k}(t)$, and $\gamma_{k},B_{k}(t)$ are defined in eqs. (48) and (49) above.
@@ C.4 Optical density matrix
$|\psi_{0}\rangle=|\Phi\rangle|\Psi\rangle$ (63)
@@
$[\rho_{c}(t)]_{nm}=a_{n}a^{*}_{m}e^{-i\delta B(t)}e^{-i\omega_{c}\Delta t}\ \mathcal{C}_{\Delta}(t)$ (65)
@@
$\mathcal{C}_{\Delta}(t)$ $=$ $\langle\Psi(t)|e^{\Delta q_{a}(\gamma_{a}a^{\dagger}-\gamma_{a}^{*}a))}e^{\Delta q_{b}(\gamma_{b}b^{\dagger}-\gamma_{b}^{*}b))}|\Psi(t)\rangle=$ (66)
$=$ $\langle\Psi|e^{\Delta q_{a}(\gamma_{a}^{*}a^{\dagger}-\gamma_{a}a))}e^{\Delta q_{b}(\gamma_{b}^{*}b^{\dagger}-\gamma_{b}b))}|\Psi\rangle$ (67)
Notice that the dependence of $\rho_{c}(t)]_{nm}$ on the initial GW state $|\Psi\rangle$ is entirely contained in $\mathcal{C}_{\Delta}(t)$, with the pre-factor being a known time-dependent function which is the same for any GW state. Therefore, if we know the initial optical state $|\Phi\rangle$ (and its decomposition in terms of the number basis) and the GW modes frequencies, by measuring the optical density matrix we can obtain information on the parent and nonlinear GW correlations. From now on we will treat $\rho_{c}(t)]_{nm}$ and $\mathcal{C}_{\Delta}(t)$ interchangeably, in the sense that knowing one implies knowledge of the other. Measuring the optical density matrix therefore allows for measurements of the GWs’ correlations. Note that we can generalize (67) to mixed states. For a general initial mixed GW state $\sigma$ (unentangled with the optical mode) we have,
$\mathcal{C}_{\Delta}(t)=\mathrm{Tr}\left(\sigma e^{\Delta q_{a}(\gamma_{a}^{*}a^{\dagger}-\gamma_{a}a))}e^{\Delta q_{b}(\gamma_{b}^{*}b^{\dagger}-\gamma_{b}b))}\right)$ (68)
@@
$\mathcal{C}_{\Delta}(t)\approx 1$ - $iq_{a}\Delta\left(\sin\omega_{a}t\langle X_{a}\rangle+(1-\cos\omega_{a}t)\langle Y_{a}\rangle\right)$ (69)
- $\frac{q_{a}^{2}\Delta^{2}}{2}\left(\sin^{2}\omega_{a}t\langle X_{a}^{2}\rangle+(1-\cos\omega_{a}t)^{2}\langle Y_{a}^{2}\rangle+\sin\omega_{a}t(1-\cos\omega_{a}t)\bigl{<}X_{a}Y_{a}+Y_{a}X_{a}\bigr{>}\right)$
- $iq_{b}\Delta\left(\sin\omega_{b}t\langle X_{b}\rangle+(1-\cos\omega_{b}t)\langle Y_{b}\rangle\right)$
- $\frac{q_{b}^{2}\Delta^{2}}{2}\left(\sin^{2}\omega_{b}t\langle X_{b}^{2}\rangle+(1-\cos\omega_{b}t)^{2}\langle Y_{b}^{2}\rangle+\sin\omega_{b}t(1-\cos\omega_{b}t)\bigl{<}X_{b}Y_{b}+Y_{b}X_{b}\bigr{>}\right)$
- $q_{a}q_{b}\Delta^{2}\big{(}\sin\omega_{a}t\sin\omega_{b}t\langle X_{a}X_{b}\rangle+\sin\omega_{a}t(1-\cos\omega_{b}t)\langle X_{a}Y_{b}\rangle$
$\ +\sin\omega_{b}t(1-\cos\omega_{a}t)\langle Y_{a}X_{b}\rangle+(1-\cos\omega_{a}t)(1-\cos\omega_{b}t)\langle Y_{a}Y_{b}\rangle\big{)}$
@@ Appendix D Parent mode quadrature in frequency doubling
$\beta(t)$ $=$ $\beta(0)+t\dot{\beta}(0)+\frac{t^{2}}{2}\ddot{\beta}(0)+$... (70)
$=$ $\beta(0)+t\left(-i\kappa\alpha^{2}(0)-\gamma_{b}\beta(0)\right)+\mathcal{O}(t^{2})$ (71)
@@
$\langle\beta(t)\rangle\approx-i\kappa t\langle\alpha^{2}(0)\rangle$ (72)
@@
$Q_{z}(t)=ze^{i\omega_{z}t}+z^{\dagger}e^{-i\omega_{z}t}\ ,\ P_{z}(t)=i\left(z^{\dagger}e^{-i\omega_{z}t}-ze^{i\omega_{z}t}\right)$ (73)
@@
$\langle Q_{a}(t)\rangle\approx\frac{\kappa t}{2}\Bigl{<}Q_{a}(0)P_{a}(0)+P_{a}(0)Q_{a}(0)\Bigr{>}$ (74)
@@
$\mathds{V}=(2\bar{n}+1)\mathbf{R}(\phi)\mathbf{S}(2r)\mathbf{R}(\phi)^{T}$ (75)
where $\mathbf{R}(\phi)$ is a $2\times 2$ rotation matrix, $\mathbf{S}(2r)=\mathrm{diag}(e^{2r},e^{-2r})$ is the squeezing matrix with parameter $r$, and $\bar{n}$ is the thermal state mean occupation number. From the definition of the covariance matrix and Eq. (75) we have,
$\frac{1}{2}\Bigl{<}XY+YX\Bigr{>}=(2\bar{n}+1)\sinh r\sin\phi\cos\phi+\langle X\rangle\langle Y\rangle$ (76)

### v2

@@ C.2 Multimode vacuum corrections and few-mode Hamiltonian
The evolution operator (C.50) contains contributions from all modes of the gravitational field, most of which are in the vacuum state. Following [13], we now show that the contribution due to empty modes can be safely neglected, as expected. The evolution of the cavity annihilation operator is given by,
$c(t)$ $=$ $\left(\prod_{k}U_{k}(t)\right)^{\dagger}c\left(\prod_{k}U_{k}(t)\right)$ (C.55)
@@
$|0\rangle=\prod_{k}|0_{k}\rangle$ (C.56)
@@
$c(t)=e^{i\mathcal{F}(t)c^{\dagger}c}e^{i\mathcal{F}(t)/2}\mathcal{G}(t)\ c$ (C.57)
@@
$\mathcal{G}(t)=\prod_{k}\langle 0_{k}|D(q_{k}\gamma_{k})|0_{k}\rangle=\exp\left[-\dfrac{1}{2}\sum_{k}q_{k}^{2}|\gamma_{k}|^{2}\right]$ (C.59)
@@
$\mathcal{F}(t)=\int\frac{d^{3}\bm{k}}{\sqrt{(2\pi)^{3}}}\,2\omega_{c}^{2}\left(\frac{8\pi G}{k^{3}}\right)\left(\omega_{k}t-\sin\omega_{k}t\right)$
$\quad\approx\left(\frac{\omega_{c}}{E_{\rm pl}}\right)^{2}\left(\int_{E_{\rm IR}}^{E_{\rm pl}}dk\right)t\approx\left(\frac{\omega_{c}}{E_{\rm pl}}\right)\omega t,$ (C.60)
@@
$\mathcal{G}(t)=\exp\!\left[-\frac{1}{2}\int\frac{d^{3}\bm{k}}{\sqrt{(2\pi)^{3}}}\,\omega_{c}^{2}\left(\frac{8\pi G}{k^{3}}\right)|\gamma_{k}|^{2}\right]$
$\approx\exp\!\left[-2\left(\frac{\omega_{c}}{E_{\rm pl}}\right)^{2}\int_{E_{\rm IR}}^{E_{\rm pl}}\frac{dk}{k}\right]$
$=\exp\!\left[-2\left(\frac{\omega_{c}}{E_{\rm pl}}\right)^{2}\ln\!\left(\frac{E_{\rm pl}}{E_{\rm IR}}\right)\right].$ (C.61)
@@
Instead of the vacuum, we could consider all modes to be populated by a thermal state at about $1\text{\,}\mathrm{K}$, the expected temperature for the cosmic GW background [86]. This would not alter $\mathcal{F}(t)$, which is state-independent, so the estimates in (C.60) remain. The $\mathcal{G}(t)$ term, however, would acquire a correction factor at most $e^{-q_{\rm pl}^{2}(2\bar{n}+1)}$, with $\bar{n}=1/(e^{\hbar\Omega_{k}/k_{B}T}-1)$ and $q_{\rm pl}$ the GW coupling strength at the Planck frequency [13]. For a GW mode of $10\text{\,}\mathrm{H}\mathrm{z}$, peak of the expected cosmic GW background spectrum, this correction factor amounts to $\approx e^{-10^{-47}}$.
@@ C.3 Few mode Hamiltonian
$H_{O}=\omega_{c}c^{\dagger}c+\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-c^{\dagger}c\left(g_{a}Q_{a}+g_{b}Q_{b}\right)$ (C.62)
@@
$P_{a}=i(a^{\dagger}-a)\ ,\ P_{b}=i(b^{\dagger}-b)$ (C.64)
in accordance do Eq. (D.77). Note the two modes couple differently to the GW detector; we have,
$\frac{q_{a}}{q_{b}}=\left(\frac{\omega_{b}}{\omega_{a}}\right)^{3/2}=2\sqrt{2}$ (C.65)
@@
$U(t)=e^{-iB(t)(c^{\dagger}c)^{2}}e^{q_{a}c^{\dagger}c(\gamma_{a}a^{\dagger}-\gamma_{a}^{*}a)}e^{q_{b}c^{\dagger}c(\gamma_{b}b^{\dagger}-\gamma_{b}^{*}b)}$ (C.66)
with $B(t)=\sum_{k=a,b}B_{k}(t)$, and $\gamma_{k},B_{k}(t)$ are defined in eqs. (C.52) and (5.13) above.
@@ C.4 Optical density matrix
$|\psi_{0}\rangle=|\Phi\rangle|\Psi\rangle$ (C.67)
@@
$\rho_{c}(t)]_{nm}=a_{n}a^{*}_{m}e^{-i\delta B(t)}e^{-i\omega_{c}\Delta t}\ \mathcal{C}_{\Delta}(t)$ (C.69)
@@
$\mathcal{C}_{\Delta}(t)$ $=$ $\langle\Psi(t)|e^{\Delta q_{a}(\gamma_{a}a^{\dagger}-\gamma_{a}^{*}a))}e^{\Delta q_{b}(\gamma_{b}b^{\dagger}-\gamma_{b}^{*}b))}|\Psi(t)\rangle=$ (C.70)
$=$ $\langle\Psi|e^{\Delta q_{a}(\gamma_{a}^{*}a^{\dagger}-\gamma_{a}a))}e^{\Delta q_{b}(\gamma_{b}^{*}b^{\dagger}-\gamma_{b}b))}|\Psi\rangle$ (C.71)
Notice that the dependence of $\rho_{c}(t)]_{nm}$ on the initial GW state $|\Psi\rangle$ is entirely contained in $\mathcal{C}_{\Delta}(t)$, with the pre-factor being a known time-dependent function which is the same for any GW state. Therefore, if we know the initial optical state $|\Phi\rangle$ (and its decomposition in terms of the number basis) and the GW modes frequencies, by measuring the optical density matrix we can obtain information on the parent and nonlinear GW correlations. From now on we will treat $\rho_{c}(t)]_{nm}$ and $\mathcal{C}_{\Delta}(t)$ interchangeably, in the sense that knowing one implies knowledge of the other. Measuring the optical density matrix therefore allows for measurements of the GWs’ correlations. Note that we can generalize (C.71) to mixed states. For a general initial mixed GW state $\sigma$ (unentangled with the optical mode) we have,
$\mathcal{C}_{\Delta}(t)=\mathrm{Tr}\left(\sigma e^{\Delta q_{a}(\gamma_{a}^{*}a^{\dagger}-\gamma_{a}a))}e^{\Delta q_{b}(\gamma_{b}^{*}b^{\dagger}-\gamma_{b}b))}\right)$ (C.72)
@@
$\mathcal{C}_{\Delta}(t)\approx 1-iq_{a}\Delta\left(\sin\omega_{a}t\langle X_{a}\rangle+(1-\cos\omega_{a}t)\langle Y_{a}\rangle\right)$
$-\frac{q_{a}^{2}\Delta^{2}}{2}\Big(\sin^{2}\omega_{a}t\,\langle X_{a}^{2}\rangle+(1-\cos\omega_{a}t)^{2}\langle Y_{a}^{2}\rangle$
$\qquad\qquad\quad+\sin\omega_{a}t(1-\cos\omega_{a}t)\langle X_{a}Y_{a}+Y_{a}X_{a}\rangle\Big)$
$-iq_{b}\Delta\left(\sin\omega_{b}t\langle X_{b}\rangle+(1-\cos\omega_{b}t)\langle Y_{b}\rangle\right)$
$-\frac{q_{b}^{2}\Delta^{2}}{2}\Big(\sin^{2}\omega_{b}t\,\langle X_{b}^{2}\rangle+(1-\cos\omega_{b}t)^{2}\langle Y_{b}^{2}\rangle$
$\qquad\qquad\quad+\sin\omega_{b}t(1-\cos\omega_{b}t)\langle X_{b}Y_{b}+Y_{b}X_{b}\rangle\Big)$
$-q_{a}q_{b}\Delta^{2}\Big(\sin\omega_{a}t\sin\omega_{b}t\langle X_{a}X_{b}\rangle+\sin\omega_{a}t(1-\cos\omega_{b}t)\langle X_{a}Y_{b}\rangle$
$\qquad\qquad\quad+\sin\omega_{b}t(1-\cos\omega_{a}t)\langle Y_{a}X_{b}\rangle+(1-\cos\omega_{a}t)(1-\cos\omega_{b}t)\langle Y_{a}Y_{b}\rangle\Big).$
@@ Appendix D Parent mode quadrature in frequency doubling
$\beta(t)$ $=$ $\beta(0)+t\dot{\beta}(0)+\frac{t^{2}}{2}\ddot{\beta}(0)+$... (D.74)
$=$ $\beta(0)+t\left(-i\kappa\alpha^{2}(0)-\gamma_{b}\beta(0)\right)+\mathcal{O}(t^{2})$ (D.75)
@@
$\langle\beta(t)\rangle\approx-i\kappa t\langle\alpha^{2}(0)\rangle$ (D.76)
@@
$Q_{z}(t)=ze^{i\omega_{z}t}+z^{\dagger}e^{-i\omega_{z}t}\ ,\ P_{z}(t)=i\left(z^{\dagger}e^{-i\omega_{z}t}-ze^{i\omega_{z}t}\right)$ (D.77)
@@
$\langle Q_{a}(t)\rangle\approx\frac{\kappa t}{2}\Bigl<Q_{a}(0)P_{a}(0)+P_{a}(0)Q_{a}(0)\Bigr$> (D.78)
@@
$\mathds{V}=(2\bar{n}+1)\mathbf{R}(\phi)\mathbf{S}(2r)\mathbf{R}(\phi)^{T}$ (D.79)
where $\mathbf{R}(\phi)$ is a $2\times 2$ rotation matrix, $\mathbf{S}(2r)=\mathrm{diag}(e^{2r},e^{-2r})$ is the squeezing matrix with parameter $r$, and $\bar{n}$ is the thermal state mean occupation number. From the definition of the covariance matrix and Eq. (D.79) we have,
$\frac{1}{2}\Bigl<XY+YX\Bigr>=(2\bar{n}+1)\sinh r\sin\phi\cos\phi+\langle X\rangle\langle Y\rangle$ (D.80)